Grade 11Organic Chemistry

Does the degree of unsaturation depend only on the number of pi bonds?Does it also differ according to different functional groups?Please give me the answer.

Vishal Kamalakannan
14 Years agoGrade 11

2 Answers

RAKSHIT  JAIN
14 Years ago

degree of unsaturation depends only on double bond and cyclic ring irrrespective of any functional group

ethanal has 1 degree of unsaturation

Jit Mitra
14 Years ago

As far as I know, degree of Unsaturation actually differs with the functional group.

 

Weightage of an element = Valency (in standard state) - 2

For,

Carbon = 4-2 = 2

Hydrogen = 1-2 = -1

Oxygen = 2-2 = 0

Nitrogen = 3-2 = 1

 

 

And Degree of Unsaturation= (sum of product (of weightage and number of corresponding element in formula)) +2)/2

 

For example,

CH3OH = (2*1+4*(-1)+0*1+2)/2 = (2-4+2)/2 = 0 (Indicates no double bond or ring)

CH3CHO = (2*2+4*(-1)+0+2)/2 = (4-4+2)/2 = 1 (Indicates one double bond or ring [and we see there is a double bond in the structure...A ring structure is also possible for the same formula (epoxide)])

 

CH3CONH2 = (2*2+5*(-1)+0+1+2)/2 = (4-5+1+2)/2 = 1 (Again a double bond and a ring is possible)..

Double bond is in amide itself. You can also try aldoxime.

For ring you can again try an epoxide.

 

Hope this is clear.